# GAR: Generative Adversarial Reinforcement Learning for Formal Theorem Proving

### v1

@@ Abstract
Solving math problems through verifiable languages such as Lean has significantly impacted both the mathematics and computer science communities. Current state-of-the-art models are often trained with expensive online Reinforcement Learning (RL) or expert iteration. However, these approaches rely on fixed problem sets, which causes inefficient training and limits the model to tackle complex problems. To overcome these limitations, we propose GAR: Generative Adversarial Reinforcement learning, a comprehensive RL training framework that jointly trains the problem composer and solver in an adversarial loop. GAR introduces an implicit curriculum learning mechanism, which aligns task difficulty with the prover’s evolving capability. It thereby improves the training efficiency and enables stronger performance of proving advanced theorems. Experiments show that with GAR training, Goedel-Prover-V2-8B and DeepSeek-Prover-V2-7B achieve an average relative improvement in pass@32 of 4.20% on MiniF2F-Test benchmark, while DeepSeek-Prover-V2’s pass@32 on ProofNet-Test increases from 22.58% to 25.81%. Beyond formal proving, GAR establishes a general RL paradigm for co-evolution of problem generation and solving under verifiable environments.
@@ 1 Introduction
Furthermore, the GAR contributes a general RL paradigm for the co-evolution of problem generation and solving in verifiable environments, which offers new directions for adversarial co-training in other reasoning-intensive domains. To facilitate the advancement of the field, we will open-source the training and inference code of GAR in the near future in here.
@@ 2.3 Summary
The GAR is an iterative framework that loops the generation stage and the RL stage details above. To provide a clear overview, we present the complete framework in the form of pseudo-code as follows:
@@ 3 Experiments
We conduct comprehensive experiments on the MiniF2F-Test (Zheng et al., 2021) and ProofNet-Test (Azerbayev et al., 2023) benchmarks to assess the performance of the GAR framework in formal proof writing. Specifically, we show in Section 3.3 that the models trained with GAR achieve better empirical results, confirm in Section 3.4 that adversarial training induces an implicit curriculum by generating progressively harder statements, and report ablation study results in Section 3.5. Due to space limitations, we have included the efficiency study and case study in Appendix B.
@@ 3.4 Statement Fuser Study
To show that GAR produces increasingly more difficult statements across iterations and thereby establishes an implicit curriculum learning, we examine the difficulty of problems generated by the statement fuser. In each iteration, we randomly sample 50 generated statements and compute their average proof correctness rate using both the base prover and GAR-trained prover at that iteration. Details of the metric we used are provided in Appendix A.2. The result of this experiment is reported in Table 3.
@@ 3.5.1 Effect of statement modification penalty
We evaluate the impact of the statement modification penalty by monitoring statement modification rates across by the provers during training rounds. Details of this metric are provided in Appendix A.3. We trained a variant of GAR on Goedel-Prover-V2-8B without the penalty in both the statement fuser and prover, and compared it with the full GAR.
@@ 5 Conclusion
Appendix A Details of the metrics
@@
A.1 Pass Rate
Pass@ is a widely used metric for evaluating formal theorem provers (Polu et al., 2022; Jiang et al., 2022; Wang et al., 2024; Xin et al., 2024a; Dong & Ma, 2025b; Wang et al., 2025b; Lin et al., 2025b; Wu et al., 2024a; Wang et al., 2025a; b). For each theorem statement, the model generates candidate proofs, which are further checked by the Lean4 verifier. If at least one of the candidates is correct, the theorem is counted as a proved theorem. The pass@ of the LLM is the fraction of theorems in the dataset with at least one correct proof generated by the prover.
A.2 Average proof correctness rate
@@
A.3 Statement modification rate
@@
Appendix B Additional experiments
@@
B.1 Efficiency Study
@@
We analyze the computational efficiency of the GAR framework, as training from scratch demands substantial resources, to evaluate this, we compare our approach with Kimina-Prover (Wang et al., 2025a), selected for its transparent reporting of training details. The number of roll-out theorems and proofs per iteration in Wang et al. (2025a) matches our configuration. From Figure 4 in their work, we observe that Kimina-Prover achieves approximately 2% of performance gains after 25 training iterations, where GAR trained Goedel-Prover-V2-8B reaches more improvements in merely three iterations. This comparison demonstrates the computational efficiency of our adversarial training paradigm. However, since current Long CoT provers have already undergone extensive RL optimization, prolonged additional RL training could degrade their performance. Given resource constraints preventing exhaustive experimentation, we leave the discussion of scalability of GAR to future works.
B.2 Case study
B.2.1 Case 1:

### v2

@@ Abstract
Solving math problems through verifiable languages such as Lean has significantly impacted both the mathematics and computer science communities. Current state-of-the-art models are often trained with expensive online Reinforcement Learning (RL) or expert iteration. However, these approaches rely on fixed problem sets, which causes inefficient training and limits the model to tackle complex problems. To overcome these limitations, we propose GAR: Generative Adversarial Reinforcement learning, a comprehensive RL training framework that jointly trains the problem composer and solver in an adversarial loop. GAR introduces an implicit curriculum learning mechanism, which aligns task difficulty with the prover’s evolving capability. It thereby improves the training efficiency and enables stronger performance of proving advanced theorems. Experiments show that with GAR training, Goedel-Prover-V2-8B and DeepSeek-Prover-V2-7B achieve an average relative improvement in pass@32 of 4.20% on MiniF2F-Test benchmark, while DeepSeek-Prover-V2’s pass@32 on ProofNet-Test increases from 22.58% to 25.81%. Beyond formal proving, GAR establishes a general RL paradigm for co-evolution of problem generation and solving under verifiable environments. The training code for this paper is open-sourced in https://github.com/RickySkywalker/GAR-Official
@@ 1 Introduction
Furthermore, the GAR contributes a general RL paradigm for the co-evolution of problem generation and solving in verifiable environments, which offers new directions for adversarial co-training in other reasoning-intensive domains. To facilitate the advancement of the field, we will open-source the training and inference code of GAR in the near future.
@@ 2.3 Summary
The GAR is an iterative framework (Wang et al., 2023) that loops the generation stage and the RL stage details above. To provide a clear overview, we present the complete framework in the form of pseudo-code as follows:
@@ 3 Experiments
We conduct comprehensive experiments on the MiniF2F-Test (Zheng et al., 2021) and ProofNet-Test (Azerbayev et al., 2023) benchmarks to assess the performance of the GAR framework in formal proof writing. Specifically, we show in Section 3.3 that the models trained with GAR achieve better empirical results, confirm in Section 3.4 that adversarial training induces an implicit curriculum by generating progressively harder statements, and report ablation study results in Section 3.5. Due to space limitations, we have included the efficiency study and case study in Appendix D.
@@ 3.4 Statement Fuser Study
To show that GAR produces increasingly more difficult statements across iterations and thereby establishes an implicit curriculum learning, we examine the difficulty of problems generated by the statement fuser. In each iteration, we randomly sample 50 generated statements and compute their average proof correctness rate using both the base prover and GAR-trained prover at that iteration. Details of the metric we used are provided in Appendix B.2. The result of this experiment is reported in Table 3.
@@ 3.5.1 Effect of statement modification penalty
We evaluate the impact of the statement modification penalty by monitoring statement modification rates across by the provers during training rounds. Details of this metric are provided in Appendix B.3. We trained a variant of GAR on Goedel-Prover-V2-8B without the penalty in both the statement fuser and prover, and compared it with the full GAR.
@@ 5 Conclusion
Ethics Statement
After carefully reviewing the ethical regulations of the conference, to the best of our knowledge, this work does not present any foreseeable ethical concerns. This research focuses exclusively on formal theorem proving using publicly available mathematical datasets, without involving human subjects, private data, or sensitive content. As far as we are concerned, no negative societal or ethical impacts are anticipated for the contribution of this work. We only use LLMs to polish the writing style and fix grammatical errors in the paper.
Reproducibility Statement
We have made efforts to ensure that our work is reproducible. The detailed description of the GAR framework, including pseudocode, data source, reward definitions, and training objectives, is presented in Section 2. The experimental setup, including benchmarks, baselines, base model choice, and hyperparameters, is presented in Section 3.1 and Section 3.2. We also plan to further ensure the reproducibility by open-sourcing the code in the near future.
Appendix A Acknowledgment
This material is based upon work supported partially by NSF under Grant No. 2416897, Grant No. 2505932, and by ORN under Grant No. N000142512318. This research used both Delta (NSF award OAC 2005572) and DeltaAI (NSF award OAC 2320345) advanced computing systems, and computing resources provided by NAIRR Pilot NAIRR250157.
Appendix B Details of the metrics
@@
B.1 Pass Rate
Pass@ is a widely used metric for evaluating formal theorem provers (Polu et al., 2022; Jiang et al., 2022; Wang et al., 2024; Xin et al., 2024a; Dong & Ma, 2025b; Wang et al., 2025b; Lin et al., 2025b; Wu et al., 2024a; Wang et al., 2025a; Li et al., 2025; Yao et al., 2025; Zhang et al., 2026). For each theorem statement, the model generates candidate proofs, which are further checked by the Lean4 verifier. If at least one of the candidates is correct, the theorem is counted as a proved theorem. The pass@ of the LLM is the fraction of theorems in the dataset with at least one correct proof generated by the prover.
B.2 Average proof correctness rate
@@
B.3 Statement modification rate
@@
Appendix C Direct comparison with Related Works
This section aims to provide a more direct comparison between GAR and other works that also try to build a dynamic dataset during training, namely Goedel-Prover-V2 (Lin et al., 2025b) and STP (Dong & Ma, 2025b).
Compared to Goedel-Prover-V2:
The statement generation in Goedel-V2 relies on a frozen large model to synthesize data for SFT. Crucially, during the RL phase, their statement set remains static. This lack of dynamic updates of statements based on the prover’s evolving capabilities may lead to suboptimal performance and efficiency as the prover outpaces the fixed problem set. In contrast, GAR continuously updates the statement fuser via adversarial RL, ensuring the generated statements remain progressively challenging and aligned with the prover’s current skill level during the RL phase, leading to better empirical performance.
Compared to STP:
While STP trains a conjecture model based on the prover’s feedback, it relies on the expert iteration framework based on offline SFT. This process is inherently inefficient because it requires generating a massive volume of training data in a single iteration to achieve effective SFT training. According to Dong & Ma (2025b), they generate 75,000 conjectures per iteration, which is larger than the entire statement set for our generation. Conversely, GAR operates within an online RL cycle. This allows the fuser and prover to continuously update more efficiently, achieving superior performance without the computational burden of the large-scale per-iteration data required by STP.
Appendix D Additional experiments
@@
D.1 Results on PutnamBench
Given that Lean4 formal reasoning is a rapidly progressing field, the MiniF2F-Test serves as a benchmark with high accuracy in current models. To further validate that the GAR training can also make the model perform better on more advanced benchmarks, we evaluate both DeepSeek-Prover-V2-7B and GAR-trained DeepSeek-Prover on PutnamBench under pass@32. The results are demonstrated in Table 5.
We can see that the GAR-trained model solves four additional problems compared to the base model on this challenging benchmark. This consistent improvement across MiniF2F, PutnamBench, and ProofNet demonstrates the robustness of the implicit curriculum established by GAR.
D.2 Efficiency Study
| Model Type | PutnamBench |
| DeepSeek-Prover-V2-7B (Ren et al., 2025) | 22/660 |
| GAR DeepSeek-Prover | 24/660 |
| Model Type | Base Model | GAR trained |
| DeepSeek-Prover-V2 (Ren et al., 2025) | 6.96% | 13.11% |
| Goedel-Prover-V2 (Lin et al., 2025b) | 24.18% | 27.05% |
@@ D.2 Efficiency Study
We analyze the computational efficiency of the GAR framework, as training from scratch demands substantial resources, to evaluate this, we compare our approach with Kimina-Prover (Wang et al., 2025a), selected for its transparent reporting of training details. The number of roll-out theorems and proofs per iteration in Wang et al. (2025a) matches our configuration. From Figure 4 in their work, we observe that Kimina-Prover achieves approximately 2% of performance gains after 25 training iterations, where GAR trained Goedel-Prover-V2-8B reaches more improvements in merely three iterations.
Notably, despite Kimina-Prover starting from an SFT base model, while GAR starts from a heavily RL-trained base model, this comparison strengthens our efficiency claim. Because it is well established in prior work (Guo et al., 2025) that models already heavily optimized via RL face diminishing returns and are inherently harder to improve than SFT baselines. The fact that GAR achieves relatively significant gains on top of RL-optimized base models in only 3-5 iterations, while Kimina-Prover gains approximately 2% from an SFT starting point over 25 iterations, demonstrates the high sample efficiency of our adversarial training paradigm. Given resource constraints that prevent exhaustive experimentation, we leave the discussion of the scalability of GAR for future work.
D.3 Statement Modification Rate Study
To dive deeper into the behavior of statement modification, we provide detailed statement modification rates of both the base model and the GAR trained model in MiniF2F-Test under pass@32. The results are shown in Table 6
From the comparison of base models, we can see that the Goedel’s increased statement modification rate happens together with its performance enhancement. We attribute this to the stronger self-correction capability it obtains. Similarly, the soft penalty in GAR ensures that the model is penalized if it simplifies the problem to a triviality. Furthermore, we can observe that when the modification rate is low, the GAR training introduces a higher modification rate, as is the case for DeepSeek-Prover. However, if the modification rate is high in the base model, the soft penalty will effectively control it within a reasonable range. These findings prove the improvement of the GAR-trained model from another point of view
D.4 Additional Ablation Studies
D.4.1 Frozen Fuser Study
To further analyze our generative adversarial training, we conduct this experiment that applies the GAR training only on the prover and keeps the statement fuser untrained. We run three iterations of Frozen fuser GAR training on Goedel-Prover-V2-8B for three iterations. The MiniF2F-Test pass@32 results are demonstrated in Table 7.
From the results, we can see that frozen fuser fails to obtain any performance improvements compared to the base model. This confirms that a static generator is unable to extend the prover’s capabilities beyond its initial limits. This proves the necessity of co-evolution for both the problem composer and the prover.
D.4.2 Single Problem Enhancement Fuiser
In order to further study the effectiveness of our idea of problem fusing, which enhances the problem difficulty by combining two problems into a single one. We conduct the experiment of replacing the statement fuser with MagicCoder (Wei et al., 2023) that enhances the problem difficulty by restating the provided single problem to make it more advanced. The results are presented in Table 7.
The result indicates that replacing the fusion process with the MagicCoder style problem-evolving strategy results in a 2.46% performance drop compared to the base model. We conclude this degradation to the limited capability of Qwen3-8B to create more advanced problems with only one reference. It may not provide sufficient difficulty gradient for the prover, leading to training failure. In contrast, our fusion approach creates a more effective implicit curriculum by combining distinct concepts from multiple problems, thereby validating the design of the Statement Fuser.
| Method | MiniF2F-Test |
| Base model | 77.87% |
| Frozen Fuser | 77.87% |
| MagicCoder Fuser | 75.41% |
| GAR trained | 80.33% |
D.5 Case study
D.5.1 Case 1: